Calculate the 1st four moment about the mean
for the following data :
x 1 2 3 4 5 6 7 8 9
y 1 6 13 25 30 22 9 5 2
1
Expert's answer
2021-07-29T13:32:41-0400

Solution:



Mean"=\mu=\dfrac{\Sigma f_i x_i}{\Sigma f_i}=\dfrac{555}{91}=6.0989\approx6.1"



Now, 1st four moments about the mean:

First order moment"=\dfrac{\Sigma f_i (x_i-\mu)}{\Sigma f_i}=\dfrac{-134.3}{91}=-1.476"

Second order moment"=\dfrac{\Sigma f_i (x_i-\mu)^2}{\Sigma f_i}=\dfrac{440.73}{91}=4.843"

Third order moment"=\dfrac{\Sigma f_i (x_i-\mu)^3}{\Sigma f_i}=\dfrac{-1115.303}{91}=-12.256"

Fourth order moment"=\dfrac{\Sigma f_i (x_i-\mu)^4}{\Sigma f_i}=\dfrac{4315.2033}{91}=47.42"